OpenStudy (anonymous):

PPPLLLEEEAAASSSEEE, help me! What is the arc length of arc CD in the circle below?

OpenStudy (anonymous):

The picture given for the question.

OpenStudy (anonymous):

@ganeshie8 @azolotor @BangkokGarrett .... Do any of you know how to do this? i know it has something to do with a formula i have, but every time I have dine that formula it's wrong.

OpenStudy (anonymous):

The length of a circular arc is given by:\[L=\frac{\alpha \pi}{180} r\] or \[L=\frac{\alpha}{360} C\]You can find the arc length by substituting in known values.

OpenStudy (anonymous):

In this case the first one will be easier since you are already given the radius.

OpenStudy (anonymous):

So would it look like this: L=15*45*3.14 / 360?

OpenStudy (anonymous):

It would be \[\frac{45\deg\times \pi}{180} \times15ft\]

OpenStudy (anonymous):

So the answer would be 11.75?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

okay cool, thanks! :)

OpenStudy (anonymous):

Can you help me with something else? @twvogels

OpenStudy (anonymous):

Yes, go ahead.

OpenStudy (anonymous):

The circle shown below has AB and BC as its tangents. If the measure of arc AC is 140°, what is the measure of angle ABC? My options are: 60° 40° 70° 50°

OpenStudy (anonymous):

The only thing that makes since is that it is 140, but thats not an option

OpenStudy (anonymous):

Is AC the arc length or the length across?

OpenStudy (anonymous):

Um... arc length i think

OpenStudy (anonymous):

Ah...I get it now. Its the angle between them. This is just a double angle problem. I was making it harder than it has to be. ABC is just \[\frac{1}{2}(220-140)=40\]

OpenStudy (anonymous):

Where did u get 220 from?

OpenStudy (anonymous):

360-140=220 If its 140deg from A to C the short way, then its 220deg the long way.

OpenStudy (anonymous):

Ok? I'm sorry. I am confused why would you need to know that there is 220 degrees everywhere else but where the triangle is covering.

OpenStudy (anonymous):

That's just the way the formula works.

OpenStudy (anonymous):

So you subtract 360 from the given degree than subtract that from the given degree again and multiply that by .5 and it will give you your answer?

OpenStudy (anonymous):

Yes, the formula is: \[m=\frac{1}{2}(major-minor)\]